OpenStudy (anonymous):

What is the value of x? -5x = -35 x=_____

OpenStudy (anonymous):

I was thinking 7 or -7?

OpenStudy (anonymous):

divide on both sides, -35 divided by -5 is positive 7

OpenStudy (anonymous):

since a negative multiplied or divided by another negative turns the answer into positive

OpenStudy (anonymous):

thank you! can you help with some more?

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

What is the value of x? x/4 - 5 = 6

OpenStudy (anonymous):

x is 2.75 add 5 on both sides, then 11 divided by 4 is 2.75

OpenStudy (anonymous):

@mathivh

OpenStudy (anonymous):

Thanks! @consetlife

OpenStudy (anonymous):

:D

OpenStudy (mathivh):

\[x/4-5=6\] Multiply both sides by 4. \[x-20=24\] Add 20 to both sides. \[x=20+24=44\]

OpenStudy (anonymous):

So its 44?